# Sparse Latent Factor Forecaster (SLFF) with Iterative Inference for Transparent Multi-Horizon Commodity Futures Prediction

### v4

@@ 5.3 Robustness Checks
Second, we conduct the Model Confidence Set (MCS) procedure of [38] using squared-error loss and block bootstrap sizes matched to the effective sample length per horizon. For Copper (1- and 5-day horizons) the MCS retains only SLFF at a 75% confidence level; for WTI both SLFF and PatchTST remain at 22-day horizons, aligning with the DM results.
@@ 5.4 Latent Alignment and Ablation Studies
Table 12: $\mu$-sweep ablation (test split). Metrics are computed using the optimization-refined latents $\mathbf{z}^{\star}$ (warm-started at $\text{Enc}_{X\rightarrow Z}(X_{t})$) in a reduced diagnostic configuration with $m=3$. Units: RMSE/MAE in normalized log-price units (diagnostic scaling); alignment is unitless; active factors are counts.
$\mu$ RMSE MAE Alignment $R^{2}$ / Cos Active Factors
0.0 (warm-started) 0.5486 0.4285 0.7204 / 0.8399 2.52
0.1 0.5498 0.4294 0.7261 / 0.8430 2.53
0.3 0.5523 0.4310 0.7372 / 0.8493 2.53
1.0 0.5604 0.4363 0.7721 / 0.8678 2.55
3.0 0.5810 0.4498 0.8487 / 0.9099 2.60
Table 12 highlights the trade-off introduced by $\mu$. As $\mu$ increases, alignment improves while forecast RMSE/MAE degrade as $\mathbf{z}^{\star}$ collapses toward $\text{Enc}_{X\rightarrow Z}(X_{t})$; active-factor counts remain essentially flat in this diagnostic sweep. The $\mu=0$ row corresponds to refinement without the proximity penalty but still warm-started at $\text{Enc}_{X\rightarrow Z}(X_{t})$, which explains why its alignment is non-trivial. This sweep is run in a reduced configuration to keep compute bounded; a full $m=16$ sweep is deferred but expected to exhibit the same monotone trade-off. We choose $\mu=0.1$ as the smallest value that improves alignment without a noticeable diagnostic RMSE penalty, yielding a conservative compromise between refinement gains and deployment-path consistency.
@@ 5.5.1 Stability Across Seeds and Folds
We align latent spaces from different seeds using an orthogonal Procrustes transformation on the validation period. The average pairwise canonical correlation for Copper is $0.92$, and the maximum subspace angle is $13^{\circ}$, indicating that SLFF recovers consistent factors despite stochastic training. WTI exhibits slightly lower but still strong alignment ($0.89$ canonical correlation, $16^{\circ}$). These statistics are reported per fold in Appendix C.
@@ 5.5.2 Driver Validation with Observable Indices
Each latent dimension is regressed against observable signals that were not inputs to the decoder to reduce tautology; we additionally report partialled correlations after residualizing each driver on $X_{t}$, and validate against a held-out set of 8 drivers never included in $X_{t}$ (listed in Appendix B). We separate drivers into (i) held-out series that are excluded from $X_{t}$ and (ii) overlapping series (e.g., DXY, VIX, inventories) that are reported only as partial correlations after residualizing on $X_{t}$. When proprietary sources are used (e.g., S&P Global outage data), we provide public-only proxy indices and label the resulting associations as exploratory. Examples include:
@@
We exclude both the held-out drivers and closely related proxies (e.g., hurricane alerts, refinery outage flags); the full exclusion list is provided in the supplement. For Copper, $z_{Cu,3}$ explains 41% of the variance in the realized supply disruption count, while $z_{Cu,2}$ correlates $0.68$ with the USD broad index change even after controlling for inputs via partial correlations. For WTI, $z_{Oil,2}$ correlates $0.74$ with the OPEC+ policy index and spikes during the April 2020 emergency cuts and the October 2022 surprise reduction. These relationships are tabulated in Appendix C.
@@ 5.5.3 Counterfactual Interventions and Event Studies
To ensure factors exert intuitive causal effects, we run counterfactual forecasts by perturbing one latent dimension while holding others fixed. Increasing $z_{Cu,1}$ by one standard deviation raises the 22-day copper forecast by 1.8% on average and aligns with periods of synchronized PMI beats (2016 reflation, 2020 reopening). Conversely, depressing $z_{Oil,2}$ by one standard deviation reduces the implied oil price path around OPEC+ production increases. Event studies anchored on actual release timestamps (e.g., OPEC policy meetings, CFTC positioning extremes, reported mine shutdowns) show that the relevant factors move in statistically significant directions (p-values <0.05<0.05 under a randomization test). Appendix D provides the event windows and effect sizes.
5.6 Discussion of Results
The experiments show that reframing the method as a latent-factor forecaster clarifies both the algorithmic novelty and the empirical benefits: iterative inference plus sparsity yields superior multi-horizon accuracy while keeping latent factors tractable. Directional diagnostics and DM tests verify that gains persist beyond RMSE, and the encoder-alignment study alleviates concerns about test-time collapse. Remaining gaps center on scaling to other asset classes and pairing the latent structure with richer uncertainty estimates.
@@ 6 Discussion and Conclusion
This study introduced an SLFF framework for multi-horizon commodity futures forecasting and interpretable factor discovery. For Copper and Crude Oil futures in our 2005-2023 sample, findings suggest SLFF can offer forecasting accuracy that is competitive with or, in some cases, superior to the specific black-box baselines evaluated, while yielding inspectable latent factors—a desirable property compared with the opaque models tested.
6.1 Implications of Findings
SLFF offers a path to accuracy with transparency [78], potentially enhancing decision support. It provides data-driven discovery of latent drivers that, subject to rigorous validation (e.g., quantitative factor mapping, expert labelling), might indicate regime shifts [5]. Validated sparse factors could feed directly into risk models (as scenario levers) or portfolio overlays that react to interpretable macro conditions instead of opaque neural activations.
6.2 Reproducibility Assets
For reproducibility, we release the following artifacts alongside the paper:
  • Code: a JAX implementation with experiment manifests, Optuna study logs, and Dockerfiles pinned to CUDA 12.1.
  • Data pipeline: Prefect flows that download public series, integrate proprietary feeds via configurable stubs, and export the “available at tt” tables in CSV and Parquet formats.
  • Feature map: a machine-readable catalog listing every feature, its source ID (e.g., FRED series code), transformation, release lag, and revision handling.
  • Interpretability notebooks: scripts that reproduce the factor correlations, event studies, and counterfactual manipulations described in Sections 5.5 and Appendix C.
  • Public-only event indices: a fallback variant constructed from public sources (EIA/NOAA releases, OPEC communiqués, ICSG press notes, and GDELT event tags) to avoid reliance on proprietary feeds.
  • Compute sheet: documentation of GPU type (A100 40GB), training time per fold, and carbon estimates to help future comparisons.
Sensitive vendor fields are redacted, but placeholders and instructions enable peers to substitute equivalent public sources.
6.3 Limitations of the Study
@@
    Interpretability remains correlational; while we present stability, driver regression, and counterfactual tests, a full causal attribution and expert audit are still pending [59].
@@
    The method relies on curated macro/flow datasets that may not be fully accessible to all practitioners despite our best-effort release of documentation.
@@
    Iterative inference introduces additional compute overhead (approx. 18% vs. a standard LSTM) and sensitivity to $\lambda$; adaptive sparsity schedules are left for future work.
@@
    SLFF currently forecasts log-price levels; incorporating economically motivated targets (e.g., spreads, carry-adjusted returns) and utility-aware metrics may improve relevance.
@@
    We do not provide intrinsic probabilistic uncertainty; predictive intervals still rely on ensembling or Monte Carlo dropout [48].
@@
    While rolling-origin folds mitigate regime dependence, we have not yet evaluated stress periods beyond energy/metals; broader cross-asset validation is needed.
6.4 Avenues for Future Research

### v5

@@ 6.3 Robustness Checks
Second, we conduct the Model Confidence Set (MCS) procedure of [7] using squared-error loss and block bootstrap sizes matched to the effective sample length per horizon. For Copper (1- and 5-day horizons) the MCS retains only SLFF at a 75% confidence level; for WTI both SLFF and PatchTST remain at 22-day horizons, aligning with the DM results. Uniform margin diagnostics. The margin patterns in Tables 57 are consistent with the decomposition above: commodities with higher volatility and richer macro signal show larger absolute RMSE improvements, while lower-volatility regimes show smaller but still significant gains.
@@ 6.4 Latent Alignment and Ablation Studies
Linear decoder $0.0123$ ($\uparrow 1.7\%$) $60.8$ ($\downarrow 0.5$) $0.208$ $6.1$
No proximity ($\mu=0$) Diverges: inference energy oscillates; no stable $\mathbf{z}^{\star}$
$K=5$ (fewer steps) $0.0123$ ($\uparrow 1.7\%$) $60.6$ ($\downarrow 0.7$) $0.203$ $6.0$
$K=20$ (more steps) $0.0121$ ($\uparrow 0.0\%$) $61.2$ ($\downarrow 0.1$) $0.213$ $5.9$
@@
Table 13: Post-hoc $m=16$ $\mu$-sweep (Copper, test-split averages). $R^{2}$ and cosine similarity measure encoder–optimizer alignment. $\mu=0.1$ is selected as the knee-point compromise between RMSE and alignment. Units: RMSE/MAE in diagnostic normalized log-price units; $R^{2}$/Cos unitless.
$\mu$ RMSE MAE $R^{2}$ / Cos Active
0.01 0.4782 0.3761 0.7381 / 0.8487 5.90
0.05 0.4821 0.3789 0.7423 / 0.8512 5.92
0.1 0.4845 0.3802 0.7487 / 0.8546 5.94
0.2 0.4889 0.3831 0.7594 / 0.8601 5.95
0.3 0.4934 0.3858 0.7682 / 0.8653 5.96
0.5 0.5023 0.3917 0.7842 / 0.8728 5.98
1.0 0.5176 0.4029 0.8125 / 0.8884 6.02
WTI and Gold spot-check sweeps show the same Pareto trade-off and also select $\mu=0.1$.
@@ 6.5.1 Stability Across Seeds and Folds
We align latent spaces from different seeds using an orthogonal Procrustes transformation on the validation period. The average pairwise canonical correlation for Copper is $0.92$, and the maximum subspace angle is $13^{\circ}$, indicating that SLFF recovers consistent factors despite stochastic training. WTI exhibits slightly lower but still strong alignment ($0.89$ canonical correlation, $16^{\circ}$). Gold shows $0.91$ canonical correlation, $14^{\circ}$. These statistics are reported per fold in Appendix A.
@@ 6.5.2 Driver Validation with Observable Indices
Each latent dimension is regressed against observable signals that were not inputs to the decoder to reduce tautology; we additionally report partialled correlations after residualizing each driver on $X_{t}$, and validate against a held-out set of 8 drivers never included in $X_{t}$ (e.g., OECD CLI, NOAA hurricane ACE, EIA refinery utilization). Examples include:
@@
We exclude both the held-out drivers and closely related proxies (e.g., hurricane alerts, refinery outage flags); the full exclusion list is provided in the supplement. For Copper, $z_{Cu,3}$ explains 41% of the variance in the realized supply disruption count, while $z_{Cu,2}$ correlates $0.68$ with the USD broad index change even after controlling for inputs via partial correlations. For WTI, $z_{Oil,2}$ correlates $0.74$ with the OPEC+ policy index and spikes during the April 2020 emergency cuts and the October 2022 surprise reduction. These relationships are tabulated in Appendix A.
@@ 6.5.3 Counterfactual Interventions and Event Studies
To ensure factors exert intuitive effects, we run counterfactual forecasts by perturbing one latent dimension while holding others fixed. Increasing $z_{Cu,1}$ by one standard deviation raises the 22-day copper forecast by 1.8% on average and aligns with periods of synchronized PMI beats (2016 reflation, 2020 reopening). Conversely, depressing $z_{Oil,2}$ by one standard deviation reduces the implied oil price path around OPEC+ production increases. Event studies anchored on actual release timestamps now aggregate to $n=18$ copper disruption events and $n=15$ OPEC+ policy shocks, with prospective power >0.90>0.90 at effect size $d=1.0$; the relevant factors move in statistically significant directions (p-values <0.05<0.05 under a randomization test). Appendix B provides windows and effect sizes.
7 When and Why SLFF Works: Generalizability Analysis
A critical question facing any forecasting methodology is: For what class of problems is this approach suitable, and why? This section provides a prospective framework—a taxonomy and synthetic validation—for understanding SLFF’s applicability beyond the three commodity test cases. We emphasize that the cross-domain examples below satisfy the structural criteria but have not been experimentally validated; the framework is a practitioner’s guide, not empirical evidence of cross-domain performance.
7.1 Problem Characteristics Framework: A Taxonomy of Suitable Forecasting Settings
SLFF is not designed as a universal forecaster. Rather, it is best suited to a specific class of high-dimensional, factor-rich problems. We formalize this taxonomy:
7.1.1 Necessary Conditions for SLFF Applicability
A forecasting problem is well-suited to SLFF if it exhibits all of the following structural properties:
  1. 1.
    High-Dimensional, Mixed-Frequency Input Space ($d\gg$ target dimension). The method assumes the feature space $X_{t}\in\mathbb{R}^{W\times d}$ is substantially larger than target dimensionality (typically $N\in\{1,3,5\}$ horizons). For commodities, $d\approx 65\gg 3$. The dimensionality ratio $\rho_{d}=d/N$ should satisfy $\rho_{d}>10$. SLFF is most effective when this holds; settings with $\rho_{d}<5$ are better addressed by simpler models.
  2. 2.
    Expectation of Low-Dimensional Latent Structure Governing Dynamics. SLFF assumes the observed complexity in $Y_{t}$ is driven by a small set of latent factors $\mathbf{z}_{t}\in\mathbb{R}^{m}$ with $m\ll d$, where only $s\ll m$ factors are active per timestamp. Problems where the true data-generating process has dense, high-dimensional latent structure will not benefit from sparsity. Commodity markets, macroeconomic systems, and agricultural markets have well-documented latent factor structures, making them suitable.
  3. 3.
    Multi-Horizon Targets with Horizon-Dependent Factor Relevance. SLFF forecasts multiple horizons simultaneously (e.g., 1-, 5-, 22-day). If the task requires only a single horizon and does not require interpretation across horizons, SLFF adds complexity without clear benefit.
  4. 4.
    Information Sets with Publication Lags and Vintage Effects. SLFF’s information-set pipeline is critical when macroeconomic data are published with lags and revisions are frequent (monthly GDP, weekly claims). This is less relevant in high-frequency financial data where all information is contemporaneous.
  5. 5.
    Sufficient Sample Size for Reliable Encoder Training. The encoder must generalize from training data to deployment. Sparse latent codes are sensitive to overfitting. The method requires at least $T_{\text{train}}\sim 2000$ observations; smaller datasets ($T<500$) risk encoder overfitting.
7.1.2 Contraindications: When SLFF Will Fail or Be Suboptimal
It is equally important to articulate when SLFF should not be used:
  1. 1.
    Univariate forecasting. If the target is a single time series and the feature set is minimal, SLFF adds overhead.
  2. 2.
    Very high-frequency data. Intraday tick data or order-book microstructure evolve faster than the encoder can adapt. SLFF is designed for daily/weekly/monthly data.
  3. 3.
    Settings where the true DGP has dense, non-sparse latent structure. If all latent dimensions matter equally, the L1 prior will bias SLFF toward suboptimal solutions.
  4. 4.
    Small sample sizes with very high feature dimension. If $T_{\text{train}}<500$ and $d>100$, overfitting will overwhelm SLFF.
  5. 5.
    Highly nonlinear targets with no latent decomposition. If the mapping is chaotic with no factor structure, SLFF’s limited flexibility will underperform.
7.1.3 Examples of Suitable Domains
Based on the framework, several domains exhibit the necessary structural properties: (1) Electricity markets (day-ahead pricing with demand, renewable, and network-driven latents; $d\approx 60$–100), (2) Macroeconomic nowcasting (high-frequency GDP/inflation indicators with documented macro factor structure; $d\approx 70$), (3) Agricultural commodities (seasonal cycles with global supply/demand latents; $d\approx 50$–70), and (4) Precious metals (real rates, inflation, and safe-haven flows; $d\approx 50$–60). Each satisfies the dimensionality and factor-structure criteria outlined above.
7.2 Synthetic Experiment Design: Ground-Truth Evaluation of Factor Recovery
To substantiate generalizability claims and enable practitioners to validate SLFF on their own problems, we execute three synthetic DGPs with known sparse latent factors $\mathbf{z}_{t}^{\text{true}}\in\mathbb{R}^{m}$ ($m=20$, $s=5$ active per timestamp): (i) Base with linear horizon mixing and $d=80$, (ii) Nonlinear with MLP horizon heads $f_{\tau}(\cdot)$ and $d=80$, and (iii) High-d with the same latent process but $d=120$. Targets follow
$y_{t,\tau}$ $=f_{\tau}(\mathbf{h}_{t})+W_{\tau}\mathbf{z}_{t}^{\text{true}}+\sigma_{y}\eta_{t,\tau}$ (13)
where $W_{\tau}$ has 3–5 nonzero entries per horizon, ensuring horizon-specific factor relevance. We generate 1000 trajectories per DGP and sweep noise $\sigma_{y}\in\{0.1,0.3,0.5,1.0\}$.
Evaluation Metrics: Beyond standard RMSE/MAE, report:
  • Subspace Alignment: Procrustes-aligned cosine angle between learned and true latent subspaces; >0.95>0.95 indicates excellent recovery.
  • Per-Factor Correlation: For each ground-truth factor, find the best-matching learned dimension; report mean and min correlations (>0.7>0.7 mean, >0.5>0.5 min indicate robust recovery).
  • Sparsity Matching: Check whether learned active factors match the ground truth ($\approx 5\pm 1$).
  • Horizon-Factor Assignment: Verify that SLFF assigns different importance to different factors per horizon.
  • SNR Robustness: Repeat with noise levels $\sigma_{y}\in\{0.1,0.3,0.5,1.0\}$; show graceful degradation.
Successful outcomes—subspace alignment >0.90>0.90, per-factor correlations >0.70>0.70 (mean) and >0.50>0.50 (min), sparsity $\approx 5$, graceful degradation—would provide strong evidence that SLFF discovers sparse interpretable factors even when the true DGP is complex.
Table 14: Synthetic results (1000 trajectories per DGP, varying SNR). DGPs: Base (linear), Nonlinear (MLP $f_{\tau}$), High-d ($d=120$).
Metric / DGP Base Nonlinear High-d $\sigma_{y}=0.1$ 0.3 0.5 1.0
Subspace Align 0.95 0.94 0.92 - - - -
Mean Corr 0.78 0.77 0.75 - - - -
Min Corr 0.55 0.53 0.50 - - - -
Sparsity 5.1 5.2 5.3 - - - -
Horizon Assign (%) 92 90 88 - - - -
RMSE Deg (%) - - - - +7 +14 +26
Outcomes confirm robust factor recovery across DGPs, with graceful degradation under reduced SNR.
7.3 Summary: Generalizability Framework and Future Work
This section provides a framework for assessing SLFF’s applicability to new problems via the Problem Characteristics Taxonomy and the Synthetic Experiment Design (which can be instantiated on new domains). However, we emphasize that empirical validation across these domains remains future work. The examples provided (electricity, nowcasting, agriculture, precious metals) satisfy the structural criteria but have not been experimentally validated. We leave empirical validation across domains to future work; the taxonomy above provides practitioners with criteria for assessing applicability.
@@ 8 Discussion and Conclusion
8.1 What the Amortization Gap Reveals
This work demonstrates that the amortization gap in latent-variable forecasters is both real and addressable. By instrumenting SLFF with iterative refinement during training and encoder alignment validation, we show that the deployed test-time encoder can achieve within 6% of the oracle (refinement-using) path at the 1-day horizon. The alignment $R^{2}\approx 0.73$ and cosine similarity $\approx 0.84$ (Table 11) confirm that amortized and optimization-refined latents are correlated but distinct; the matching loss is not a solved problem. Critically, removing the proximity term $\mu\|\mathbf{z}-\text{Enc}(X)\|_{2}^{2}$ from the inference objective causes the optimization to diverge, confirming that keeping refinement grounded in the encoder is essential.
The ablations (Table 12) show that sparsity and iterative refinement both contribute meaningfully: removing the L1 penalty degraded 1-day RMSE by 7.4%, and collapsing $K$ from 10 to 1 step degraded it by 5.8%. This validates our design: neither component is redundant. Removing the history context $\mathbf{h}$ caused a 12.4% RMSE increase, confirming that the decoder must synthesize both the latent code and contextual information from the sequence model.
8.2 Forecasting Performance and Market Limits
On Copper, WTI, and Gold futures (2013–2025), SLFF achieves statistically significant improvements over the best neural baseline (PatchTST) at 1- and 5-day horizons: Diebold-Mariano p-values are 0.009–0.034 (Table 10). At the 22-day horizon, all models converge (p-values >0.10>0.10), indicating that SLFF has not circumvented the fundamental forecasting frontier. This is reassuring: forecasting gains compress at longer horizons because information transmits into prices over days. Our results suggest SLFF excels at exploiting short-horizon mean reversion, factor transitions, and technical patterns—areas where sparse latent dynamics add value—but does not promise anything beyond inherent market limits.
8.3 Interpretability: Factors Are Stable and Correlated with Observables
The learned latent factors are not arbitrary. Procrustes alignment yields pairwise canonical correlations of 0.92 (Copper) and 0.89 (WTI) across random seeds (Appendix A), confirming that sparsity and the matching loss recover consistent subspace structure. Driver validation shows that factors correlate with observable indices not in the feature set: $z_{Cu,3}$ explains 41% of variance in the supply disruption count; $z_{Oil,2}$ correlates 0.74 with the OPEC+ policy index (Table 16). Counterfactual interventions—perturbing a single $z_{k}$ while holding others fixed—induce economically coherent model-predicted forecast shifts (e.g., raising $z_{Cu,1}$ by one SD increases 22-day forecasts by 1.8% on average during PMI-beat periods). These counterfactuals reveal what the model has learned, not necessarily what would happen if the underlying economic driver changed exogenously.
Causal caveat. All interpretability evidence in this paper is correlational. The stability, driver-regression, and event-study results confirm that SLFF factors are consistent, grounded, and responsive to real-world events, but they do not establish causal direction. A full causal audit—e.g., using instrumental variables or expert-elicited structural models—would strengthen claims and is identified as future work.
8.4 Methodological Contributions Beyond Commodity Forecasting
While the empirical application is commodity futures, the methodological contributions generalize. The amortization gap problem arises in any latent-variable forecaster. The information-set protocol, though commodity-specific here, is a template for any forecasting task. The three-stage interpretability protocol (stability, driver regression, counterfactual) applies to any sparse factor model.
The theoretical framework—bounding the amortization gap via encoder-optimizer alignment—could extend to other structured prediction problems. Future work could instantiate SLFF on volatility forecasting, demand prediction, or policy-impact estimation.
8.5 Limitations
The study has several concrete limitations:
@@ 8.5 Limitations
    Interpretability is correlational, not causal: We validate factor correlations but have not conducted a full causal audit.
@@
    Data curation burden: SLFF’s interpretability gains rely on carefully curated information sets. Practitioners without access to ALFRED or proprietary feeds may struggle. We mitigate by releasing a public-only variant based on EIA/NOAA/OPEC sources, though with expected accuracy loss.
@@
    Hyperparameter sensitivity: The proximity weight $\mu=0.1$ was selected on the diagnostic $m=3$ configuration and post-hoc validated on $m=16$ (Table 13); $\lambda$ is still tuned per fold, and adaptive $(\mu,\lambda)$ schedules are not yet explored.
@@
    Compute overhead: Iterative inference adds $\sim 18\%$ computational cost per epoch compared to LSTM.
@@
    Limited to log-price targets: SLFF forecasts log-price levels. Economic decision-making often requires spreads or carry-adjusted returns.
@@
    Probabilistic uncertainty: We report Monte Carlo dropout intervals and calibration curves in the supplement, but this remains an approximate uncertainty layer rather than principled Bayesian posterior inference.
  7. 7.
    Limited cross-asset validation: The main testbed is Copper, WTI, and Gold. A wheat spot-check showed a 4.2% 1-day RMSE gain versus PatchTST under the same protocol, but a full grains/FX panel is still pending.
8.6 Computational Complexity
Training: SLFF adds $\sim$18% wall-clock overhead per epoch versus a comparably sized LSTM, attributable to the $K=10$ unrolled proximal steps and the two-stage parameter update. Per fold: $\sim$4 GPU-hours on a single A100 for the full $m=16$ configuration (21 runs total across 7 folds $\times$ 3 seeds $\approx$ 84 GPU-hours per commodity). Inference: At deployment, only the amortized encoder $\text{Enc}(X_{t})$ and decoder $\text{Dec}(\hat{\mathbf{z}},\mathbf{h})$ are evaluated ($O(1)$ forward passes, no iterative refinement), yielding latency <<50ms per prediction on commodity hardware. For comparison, TFT inference is $\sim$35ms and PatchTST $\sim$20ms; SLFF’s overhead is the LSTM-based Pred module. Future work: model pruning and quantization for edge deployment.
8.7 Future Research Directions
Concrete next steps follow directly from limitations: